The Ka for benzoic acid is 6.5 x 10^-5.

Calculate the pH of a 0.15 M benzoic acid solution.


1
Expert's answer
2020-12-01T09:05:21-0500

The formula for Ka is:

Ka = [H+][B-]/[HB]

where:

[H+] = concentration of H+ ions

[B-] = concentration of conjugate base ions

[HB] = concentration of undissociated acid molecules

for a reaction HB → H+ + B-

Benzoic acid dissociates one H+ ion for every C6H5COO- ion, so [H+] = [C6H5COO-].

Let x represent the concentration of H+ that dissociates from HB, then [HB] = C - x where C is the initial concentration.

Enter these values into the Ka equation:

Ka = x · x / (C -x)

Ka = x²/(C - x)

(C - x)Ka = x²

x² = CKa - xKa

x² + Kax - CKa = 0

Solve for x using the quadratic equation:

x = [-b ± (b² - 4ac)½]/2a

x = [-Ka + (Ka² + 4CKa)½]/2

Ka = 6.5 x 10-5

C = 0.15 M

x = {-6.5 x 10-5 + [(6.5 x 10-5)² + 4(0.15)(6.5 x 10-5)]½}/2

x = (-6.5 x 10-5 + 1.6 x 10-3)/2

x = (1.5 x 10-3)/2

x = 7.7 x 10-4

Find pH:

pH = -log[H+]

pH = -log(x)

pH = -log(7.7 x 10-4)

pH = -(-3.11)

pH = 3.11